a motorboat travels 168 km in 3 hours going upstream and 800 Km in 8 hours going downstream. What is the rate of the boat in still water and what is the rate of the current?
Answer by Alan3354(69443)

a motorboat travels 168 km in 3 hours going upstream and 800 Km in 8 hours going downstream. What is the rate of the boat in still water and what is the rate of the current?
--------------
168/3 = 56 kph
800/8 = 100 kph
The boat's speed is the average = 78 kph
The current is the difference = 78 - 56 = 22 kph